Circular No. 7145

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999bu IN NGC 3786 W. D. Li, University of California at Berkeley, on behalf of the Lick

Observatory Supernova Search (cf. IAUC 6627, 7126) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT), reports the discovery of an apparent supernova in NGC 3786, a Seyfert galaxy that is in a pair of interacting galaxies (Arp 294 and NGC 3788). SN 1999bu was discovered on an unfiltered image taken on Apr. 16.3 UT (mag ∼ 17.5). Inspection of earlier images taken by KAIT reveals that the supernova was barely visible on Mar. 13.3 and clearly present on Mar. 28.3 (mag ∼ 17.6). The new object is located at α = 11^{h}39^{m}42^{s}.50, δ = +31^{o}54′29′′.1 (equinox 2000.0), which is ∼ 1′′.3 west and 3′′.1 south of the nucleus of NGC 3786. A KAIT image of the same field on Mar. 7.3 (limiting mag ∼ 18.5) shows nothing at the position of the supernova.

COMET 1999 H1 The following additional precise CCD positions are available:

1999 UT α_{2000} δ_{2000} Observer
Apr. 16.77635 13^{h}46^{m}26.^{s}62 −68^{o}20′21′′.9 Garradd
16.77679 13 46 26.17 −68 20 21.8  _{”}
16.77727 13 46 25.66 −68 20 21.6  _{”}
17.65347 13 31 54.34 −68 11 39.8 Zoltowski
17.65748 13 31 50.22 −68 11 36.5  _{”}
17.66308 13 31 44.56 −68 11 32.1  _{”}
17.66996 13 31 37.60 −68 11 26.6  _{”}

G. J. Garradd (Loomberah). 0.45-m f/5.4 reflector. m_{2} = 13.9. F. B. Zoltowski (Woomera). 0.30-m f/10 Schmidt-Cassegrain reflector.

Visual m_{1} and coma-diameter estimates: Apr. 17.31 UT, 9.6, 2′.2 (S. T. Rae, Palmerston North, New Zealand, 0.25-m reflector; comet well condensed); Apr. 17.46, 9.2, 2′.0 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 17.67, 8.9, 3′ (Rae, 0.11-m reflector).

COMET P/1998 U3 (JÄGER) Visual m_{1} estimates: Mar. 4.75 UT, 10.4 (M. Lehký, Hradec Králové,

Czech Rep., 25×100 binoculars); 17.79, 11.1 (W. Hasubick, Buchloe, Germany, 0.44-m reflector); Apr. 6.05, 11.1 (J. Bortle, Stormville, NY, 0.41-m reflector); 15.88, 11.6 (A. Baransky, Kiev, Ukraine, 0.20-m reflector).

1999 April 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green